PLEASE SHOW ALL WORK

As you can see, i got the first part wrong :/

1. A particular accelerator can accelerate electrons
Answer #1

a) lemda = h/p = h / sqrt(2mE) = 6.62*10^-34/ sqrt(2*3.1*10^-31*37*10^9*1.6*10^-19) = 1.10*10^-14 m

b) d/lemda = 2/11 = 0.18 fraction of proton diameter can such particle probe